OpenStudy (korosh23):

Physics question:

OpenStudy (korosh23):

When do we use m1v1= m2v2 ??? when do we use m1v1 + m2v2= M1V1+ M2V2 --> Conservation of momentum, when momentum is conserved.

OpenStudy (korosh23):

When do we use those formula?

OpenStudy (michele_laino):

the second formula can be used when we have an interaction between two particles, and the system composed by those two particles is isolated, namely no external forces are acting on each particle, only internal forces

OpenStudy (korosh23):

ok

OpenStudy (michele_laino):

the first formula, can be used when we have to study a collision between two particles, say particle #1 and particle #2. For example when are considering the bullet fired by a gun. There, the momentum gained by the bullet and the one gained by the gun are opposite with the same direction, and their magnitudes checks this condition: m_g*v_g=m_b*v_b where b stands for bullet, and g stands for gun

OpenStudy (michele_laino):

again neglecting any external force acting on both bullet and gun

OpenStudy (korosh23):

So we can use formula 1 only when there is collision?

OpenStudy (michele_laino):

more precisely when we have a particular collision, where both particle have the same mass, furthermore, before collision, particle #2 is at the rest. Otherwise, we can use the first formula when we have to study the interaction bullet-gun, for example

OpenStudy (korosh23):

Oh I got it

OpenStudy (korosh23):

Formula #1 is when collisions happen and there are two objects. They do not add their mass together Formula #2 when collisions happen and momentum conserved, so masses add together and we have to find the velocities. MV1+MV2= (M1+M2)v

OpenStudy (michele_laino):

if we add mass together, after collision, then that collision is not elastic for example

OpenStudy (michele_laino):

the last formula: MV1+MV2= (M1+M2)v can be used to study, for example an inelastic collision

OpenStudy (korosh23):

Oh ok. That makes sense, Thank you for your helpful explanations. You are such a good tutor.

OpenStudy (michele_laino):

:)